x=cos

3

t,y=sin

3

t,t∈[0,2π]

x=cos3⁡t,y=sin3⁡t,t∈[0,2π]




1
Expert's answer
2021-09-14T00:01:27-0400
"x=\cos(3t), y=\sin(3t), t\in[0,2\pi]."

The length of the parametric curve is given by


"L=\displaystyle\int_{0}^{2\pi}\sqrt{(x'(t))^2+(y'(t))^2}dt"

"=\displaystyle\int_{0}^{2\pi}\sqrt{(-3\sin(3t))^2+(3\cos(3t))^2}dt"

"=\displaystyle\int_{0}^{2\pi}3dt=[3t]\begin{matrix}
 2\pi \\
 0
\end{matrix}=6\pi (units)"



The length of curve is "6\pi."